OpenStudy (anonymous):

which expression is equal to (f+g)(x) f(x)=x(exponent2) + 10;g(x) = x - 3 A) 2x(exponent2) +7 B)x(exponent2) + x + 7 C) x(exponent2) + x + 13 D) 7 x(exponent2) + x - 3

OpenStudy (jango_in_dtown):

(f+g)(x)=f(x)+g(x)=x(exponent2) +10+x - 3=x(exponent2) + x + 7